# Carve3D: Improving Multi-view Reconstruction Consistency for Diffusion Models with RL Finetuning

### 21 Dec 2023 v1

@@ 5 Experiments
In this section, knowing that Carve3D’s RLFT steadily improves the MRC reward on the training set (Figure 5), we aim to answer the following questions:
@@
    Does Carve3D’s improved MRC generalize to the testing set?
@@
    Qualitatively, is Carve3D more consistent than the base model? And comparing to longer SFT?
@@
    Qualitatively, does Carve3D sacrifice the diversity, prompt alignment, and texture details of the base model?
We quantitatively and qualitatively compare Carve3D with three versions of Instant3D [24], with 10K, 20K, and 100K SFT steps respectively. Instant3D-10K is both the default model in [24] and also our base model. For fair qualitative comparisons, results from each model are generated from the the same initial noise. Since 20K and 100K versions of Instant3D and Carve3D are all finetuned from Instant3D-10K, their output tend to represent the same object when given the same initial noise (e.g. Figures 1 and 6).
5.1 Comparison with Base Model and Longer SFT
@@
Carve3D (Ours) 0.0606
Table 1: Carve3D (RLFT with MRC reward on Instant3D-10k), achieves substantially better MRC than baselines, which corresponds to better multi-view 3D consistency. We evaluate these text-to-multiview diffusion models on the DreamFusion testing set, containing 414 text prompts. We generate 4 outputs for each prompt, and the average MRC is computed over the 1656 outputs. For each output, we use the same randomly sampled initial noise for all models to ensure the comparison is fair.
@@
As shown in Table 1, when evaluated on the testing set, Carve3D achieves substantially improved MRC over the base model. More SFT steps indeed provides better multi-view consistency and achieves better MRC, with Instant3D’s 100K version performing the best and 10K version performing the worst. However, Carve3D still outperforms even the most consistent 100K version of Instant3D by a notable gap. This suggests that the explicit multi-view consistency objective in MRC, paired with our RLFT algorithm, can improve the model’s consistency more efficiently than SFT. Furthermore, our RLFT provides generalization of the improved multi-view consistency, although only finetuned on 30 prompts. Such generalization, also observed in [5, 36], is likely derived from the strong knowledge from the base model.
@@
Figure 12 shows the improved multi-view consistency and the resulting NeRF reconstruction quality. While the multi-view images generated by the base model may be inconsistent, causing artifacts such as floaters and broken geometry, Carve3D can fix such inconsistency in the multi-view images and produce NeRF with clean geometry, free of artifacts. In Figure 11, Carve3D continues to show superior multi-view consistency and reduced NeRF artifacts, but such improvement is less and further less obvious when compared to the 20K and 100K version of Instant3D [24], which aligns with the qualitative results in Table 1.
@@
By virtue of our RLFT with KL-divergence regularization (Section 4.2), which prevents distribution shift, and curated low-reward prompts, which describes complex objects (Section C.1), Carve3D preserves the prompt alignment and the texture details of the base model, as shown in Figure 12. On the other hand, longer SFT causes additional distribution shift in Instant3D [24] from the base SDXL [39] towards the SFT training set [12]. As shown in Figure 11, Instant3D-20K and Instant3D-100K exhibits degradation in diversity, realism, and level of detailed textures. This quality degradation with longer SFT is also observed in [24].
@@
As shown in Figure 13, Carve3D can preserve the diversity of the base model. This owes to our RLFT process, which repeatedly samples different initial noises for the diffusion model to generate diverse results (Figure 2).
5.2 Evaluating Existing Methods
At the time of writing, MVDream [46] is the only text-to-multi-view diffusion model with released code and model weights. As shown in Tables 1 and 11, its outputs have notably worse multi-view consistency, realism, and level of details than all three Instant3D variants and Carve3D. This supports our claim in Section 1 that current multi-view diffusion models that solely rely on SFT still suffer from the inconsistency problem and could benefit from RL finetuning.
@@ 5.3 User Study
In addition to the quantitative and qualitative comparisons in Section 5.1, we conducted a user study to further understand the qualitative results of Carve3D’s RLFT when perceived by human. To run the study we randomly selected 20 unseen testing prompts. For each text prompt, we generated a pair of data from both the base and the finetuned models with the same initial noise. Then, we provided both the tiled 4-view image and the turntable video of the reconstructed NeRF to participants and asked them the following two questions: (1) Which result is more 3D-consistent? and (2) Which result is better aligned with the prompt? As shown in Figure 8, 68.33% of participants believe that Carve3D’s generated results are more 3D consistent than that of the base model [24]. Given that the multi-view consistency in the base model has already been much improved with SFT  [^2] Please see https://jiahao.ai/instant3d/ for base model’s 3D consistency, the nearly $37\%$ gain in human preference introduced by Carve3D on randomly selected testing prompts is impressive. Furthermore, the Carve3D finetuned model exhibits similar prompt alignment, as participants’ votes are evenly distributed among ours, base model, and “same”. The preservation of alignment can be attributed to the KL divergence regularization (Section 4.2) as well as early stopping the RLFT according to KL divergence (Section 4.2.3).
@@ 6 Conclusion
In this paper, we propose Carve3D, an RL finetuning method to improve the reconstruction consistency of 2D diffusion models. The reward of Carve3D relies on MRC, a novel metric that measures the reconstruction consistency by comparing input multi-view images with the renderings of the reconstructed NeRF at the same viewpoints. The effectiveness and robustness of MRC are also validated by showing its correlation with intentional distortions. Lastly, we conduct experiments and a user study to show that Carve3D significantly improves the reconstruction consistency of multi-view images and the resulting quality of the NeRF. These enhancements are achieved without sacrificing the prompt alignment, texture details, or prompt alignment of the base model.
Our MRC metric can serve as a valuable tool for evaluating any multi-view generative methods and guiding future developments in the field. Although we only demonstrate our RLFT with MRC on one multi-view diffusion model [24], it can be directly adapted to other text-to-multi-view diffusion models; such adaptation only requires tuning a few hyperparameters related to the scaling laws for diffusion model RLFT (Section 4.2.3). Our surprising finding behind the choice of REINFORCE [53] over PPO [45] for better training stability could also be applied in broader RLFT scenarios.
As AI models grow more powerful, it becomes more important to evaluate and improve their safety and reduce their bias. RLFT has been widely used for LLM alignment as it allows models to be finetuned with hard-to-specify objectives and its results are generalizable without undermining the base model’s knowledge. As the first work to use RLFT for text-to-3D and on diffusion models at the SDXL scale, we hope Carve3D can inspire more alignment research in the computer vision community.
@@ Appendix A Appendix Summary
In the appendix, we cover additional details of DDPO and RLFT (Appendix B, the training data details (Section C.1), training details (Section C.2), additional MRC metric experiments (Appendix D), ablation studies (Appendix E), and future work (Appendix F).
@@ B.1 Definitions
Following [45, 36, 54, 5], an epoch is defined as one round of data collection (sampling), which may consists multiple PPO update steps (training), as discussed in Equations 6 and 4.2. This definition of “epoch” is different from the meaning in supervised learning which usually refers to go through all data once. Since we opt for using pure on-policy training (vanilla policy gradient), as discussed in Section 4.2, we only do one training step per sampling step, and thus our sampling batch size and training batch size are equal.
@@ B.2 $\text{DDPO}_{\text{IS}}$ Policy Gradient Function
By using the advantage term (Equation 1) in place of the reward, the $\text{DDPO}_{\text{IS}}$ policy gradient function is:
@@ B.3 Hypotheses on Stability and Sample Efficiency
Diverging from DDPO [5] and most LLM RLHF literature [36, 54, 50, 2, 3], we choose REINFORCE [53] ($\text{DDPO}_{\text{SF}}$) over PPO [45] ($\text{DDPO}_{\text{IS}}$) for its superior training stability. We provide two hypotheses behind our surprising finding.
@@
(2) The RLFT sample efficiency is less important for a large model which requires less finetuning steps, as demonstrated in studies of LLM instruction finetuning [10]. Similarly, our RLFT on a 2.6B-parameter UNet from SDXL [39] only takes 55 epochs, as opposed to DDPO’s [5] RLFT on a 860M-parameter UNet from SD 1.4 [42] using 200 epochs. Therefore, the potential sample efficiency gain provided by the multi-step update of PPO [45] gets outweighted by the training stability provided by REINFORCE [53].
@@ C.1 Training Data
An advantage of RLFT over SFT is that, we can manually create a high-quality text prompts training set, while creating a dataset of diverse ground truth multi-view images for these high-quality text prompts is prohibitively expensive for SFT. By relying on samples generated by the model itself to compute the reward and the loss, RLFT can optimize a model beyond the limitation of a dataset and preserves the diversity and the style of the base model. In Carve3D, our training prompts preparation process involves two strategies.
@@
Instead of randomly sampling prompts from a dataset, we employ a data curation strategy where prompts with lowest rewards are selected. Specifically, we run inference of the base model on a prompt dataset, generating four results per prompt, compute the MRC rewards for each result, and sort the prompts according to their average reward. This is derived from observation that, for certain prompts, the model generates nearly optimal outputs with rewards close to the rewards of ground truth views of a 3D asset [12] (Figure 4). Thus, the curated lowest-reward prompts have substantial room for 3D consistency improvement and prevent learning stagnation. This approach not only brings more efficient training but also provides a more generalized improvement in 3D consistency to the testing set.
@@
The prompt dataset from DreamFusion [40], which contains 414 prompts and is commonly used as testing set. To employ the DreamFusion prompt set also as our testing set, we create a new prompt dataset with ChatGPT4 [35]. Following our training data curation strategy, we first sort the DreamFusion [40] prompts according to their rewards attained by the base Instant3D [24] model. We provide the sorted prompt set to ChatGPT4, and ask it to summarize the characteristics of the low-reward prompts by looking at the low-reward, median-reward, and high-reward prompts. ChatGPT4 summarizes low-reward prompts to possess properties of “complex and creative”. We then ask it to generate 100 low-reward prompts that are both complex and creative, and another 100 low-reward prompts that are “complex but not too creative”. For each set, again, we sort the prompts according to their rewards, and select those with the lowest rewards to be our training prompt set. Our best results are obtained with the “complex but not too creative” set.
@@ C.2 Training Details
We use minibatches of size 8 during sampling and 4 during training due to the limited GPU memory. The total batch size of 768 is evenly distributed among each GPU, so that the per GPU batch size is 16. The model samples two minibatches of size 8 on all GPUs to reach the total batch size. Similarly, the model accumulates gradients over four minibatches of size 4 on all GPUs, before synchronizing the gradients and performing an optimizer step. We use a per prompt stat tracker with windows of size 76, so that it roughly tracks all the rewards per prompt ever 3 epochs. This is much larger than DDPO’s default tracking window of size 32 for better training stability. The coefficients for the advantage terms in Equation 5 are $\alpha=1$ and $\beta=0.2$.
@@ Appendix D Additional MRC Metric Experiments
Here, we show the full results for the metric experiments for the inpainting distortion (Figure 14) discussed in Sections 3.3, 3 and 4. We also conduct metric experiments with other distortions types: azimuth rotation (Figure 15, and elevation rotation (Figure 16). In azimuth and elevation rotation, for one out of the four views, we rotate the object with an azimuth or elevation rotation by $3.6$ or $4$ degrees, before rendering that view, and also use the original camera extrinsic matrix as the input to Sparse View LRM. The quantitative results matches our expectations, where MRC (LPIPS) monotonically decreases as we intentionally add more distortion.
@@
Here, we compare substituting LPIPS with L1, L2, PSNR, and SSIM in the metric experiments on all distortion types. In the inpainiting distortion experiments (Figure 14), which is the most representative of diffusion model’s inconsistencies, LPIPS is more linear than other pixel level image metrics. In azimuth and elevation distortion experiments (Figures 15 and 16), all image metrics shows monotonically decreasing pattern, while pixel-level image metrics are more linear. This is expected as the distortion is pixel-aligned and more structured.
Figure 9: Alation study on the boundingbox normalization of LPIPS for MRC evaluation. Top: with normalization, the 3D objects keep similar size after finetuning. Bottom: without normalization, the sizes of 3D objects are greatly reduced after RL finetuning.
Figure 10: Ablation study on KL divergence regularization. Top: KL Divergence between the base model and the finetuned model on testing set. Bottom: mean MRC reward on testing set. Our KL divergence regularization does not sacrifice the model’s efficiency on improving the reward. Without KL divergence regularization, the finetuned model’s KL divergence to the base model on the training and the testing set grows faster, which results in degraded object identity and reduced texture details.

### 09 Apr 2024 v2

@@ 5 Experiments
In this section, we evaluate the Carve3D Model (Carve3DM), obtained by applying the Carve3D Reinforcement Learning Finetuning (RLFT) algorithm on Instant3D [24]. See Sec. C.2 for our training details, Sec. C.3 for our experiment settings and Appendix E for our ablation studies.
5.1 Comparison with Base Model and Longer SFT
As shown in Fig. 5, Carve3DM’s Multi-view Reconstruction Consistency (MRC) reward steadily improves on the training set. We aim to further answer the following questions:
@@ 5.1 Comparison with Base Model and Longer SFT
    Does Carve3DM’s improved MRC generalize to the testing set?
@@
    Quantitatively, is Carve3DM more consistent than the base model and the model with more Supervised Finetuning (SFT) steps?
@@
    Qualitatively, does Carve3DM sacrifice the diversity, prompt alignment, and texture details of the base model?
We quantitatively and qualitatively compare Carve3DM with Instant3D models with 10K, 20K, and 100K SFT steps, where Instant3D-10K is the default model in [24].
@@
Carve3DM (Ours) 0.0606
Zero123++ 0.0700
SyncDreamer 0.1018
Table 1: Quantitative comparison of MVDream [46], Instant3D [24] with 10K (the base model), 20K, and 100K SFT steps, Carve3DM (ours, finetuned from Instant3D-10K), Zero123++ [47], and SyncDreamer [28]. We evaluate them by generating 4 outputs for each of the 414 text prompts in the DreamFusion [40] testing set. We let Zero123++ and SyncDreamer to use one of Carve3DM’s output multi-view images as their input image conditioning. Carve3DM achieves substantially better MRC than all baselines, indicating its superior multi-view consistency.
@@
As shown in Tab. 1, when evaluated on the testing set, Carve3D achieves substantially improved MRC over the base model. More SFT steps indeed provides better multi-view consistency and achieves better MRC, i.e. Instant3D’s 100K version is better than the 20K version, which is better than the 10K version. However, Carve3DM still outperforms even the most consistent 100K version of Instant3D by a notable gap. This suggests that the explicit multi-view consistency objective in MRC, paired with our RLFT algorithm, can improve the model’s consistency more efficiently than SFT.
Furthermore, although our RLFT only uses 30 training prompts, it brings multi-view consistency improvement that generalizes to the testing set containing 415 prompts. Such generalization, also observed in [5, 36], is likely derived from the strong knowledge from the base model.
@@
Appendix Fig. 12 shows the improved multi-view consistency and the resulting Neural Radiance Field (NeRF) reconstruction quality. While the multi-view images generated by the base model may be inconsistent, causing artifacts such as floaters and broken geometry, Carve3D can fix such inconsistency in the multi-view images and produce NeRF with clean geometry, free of artifacts. In Appendix Fig. 11, Carve3DM continues to show superior multi-view consistency and reduced NeRF artifacts, but such improvement is less obvious when compared to the 20K and 100K version of Instant3D [24], similar to our quantitative results in Tab. 1.
@@
By virtue of our RLFT with KL-divergence regularization (Sec. 4.2.2), which prevents distribution shift, and our curation of training prompt dataset (Sec. C.1), Carve3DM preserves the prompt alignment and the texture details of the base model, as shown in Appendix Fig. 12. On the other hand, longer SFT causes additional distribution shift in Instant3D [24] from the base SDXL [39] towards the SFT training set [12]. As shown in Appendix Fig. 11, Instant3D-20K and Instant3D-100K exhibits degradation in diversity, realism, and level of detailed textures. This quality degradation with longer SFT is also observed in [24].
@@
As shown in Appendix Fig. 13, Carve3DM preserves the generation diversity of the base model. This owes to our RLFT process, which repeatedly samples different initial noises for the diffusion model to generate diverse results (Fig. 2).
5.2 Comparison with Existing Methods
We further compare Carve3DM with the text-conditioned multi-view diffusion model, MVDream [46], and two image-conditioned models, Zero123++ [47] and SyncDreamer [28]. As shown in Tab. 1 and Appendix Fig. 11, Carve3DM’s outputs have notably better multi-view consistency, realism, and level of details than the three baselines.
@@ 5.3 User Study
In addition to the quantitative and qualitative comparisons in Sec. 5.1, we conduct a user study to further understand the qualitative results of Carve3DM when perceived by human. To run the user study, we randomly selected 20 unseen testing prompts. For each text prompt, we generated a pair of data from both the base and the finetuned models with the same initial noise. Then, we provided both the tiled 4-view image and the turntable video of the reconstructed NeRF to participants and asked them the following two questions: (1) Which result is more 3D-consistent? and (2) Which result is better aligned with the prompt? As shown in Fig. 8, 68.33% of participants believe that Carve3DM’s generated results are more 3D consistent than that of the base model [24]. Given that the multi-view consistency in the base model has already been much improved with SFT  [^2] Please see https://jiahao.ai/instant3d/ for base model’s 3D consistency, the nearly $37\%$ gain in human preference introduced by Carve3D on randomly selected testing prompts is impressive. Furthermore, participants find that Carve3DM exhibits similar prompt alignment as Instant3D. The preservation of alignment can be attributed to the KL divergence regularization (Sec. 4.2.2) and early stopping the RLFT regarding KL divergence (Sec. 4.2.3).
@@ 6 Conclusion
In this paper, we propose Carve3D, a Reinforcement Learning Finetuning algorithm to improve the reconstruction consistency of 2D multi-view diffusion models. Carve3D relies on MRC, a novel metric that measures the reconstruction consistency by comparing multi-view images with their corresponding NeRF renderings at the same viewpoints. The resulting model, Carve3DM, demonstrates substantially improved multi-view consistency and NeRF quality without sacrificing the prompt alignment, texture details, or prompt alignment of the base model. Our results conclude that pairing SFT with Carve3D’s RLFT is essential for developing multi-view-consistent diffusion models.
@@ Appendix A Appendix Summary
In the appendix, we cover additional details of DDPO [5] and Reinforcement Learning Finetuning (RLFT) (Appendix B), training data details (Sec. C.1), training details (Sec. C.2), experiment details (Sec. C.3), additional Multi-view Reconstruction Consistency (MRC) metric experiments (Appendix D), ablation studies (Appendix E), and broader impacts and future work (Appendix F).
@@ B.1 Definitions
Following [45, 36, 55, 5], an epoch is defined as one round of data collection (sampling), which may consists multiple PPO [45] update steps (training), as discussed in Eqs. 11 and 4.2. This definition of “epoch” is different from the meaning in supervised learning which usually refers to go through all data once. Since we opt for using pure on-policy training (vanilla policy gradient), as discussed in Sec. 4.2, we only do one training step per sampling step, and thus our sampling batch size and training batch size are equal.
@@ B.2 $\text{DDPO}_{\text{IS}}$ Policy Gradient Function
By using the advantage term (Eq. 6) in place of the reward, the $\text{DDPO}_{\text{IS}}$ policy gradient function is:
@@ B.3 Hypotheses on Stability and Sample Efficiency
Diverging from DDPO [5] and most Large Language Model (LLM) Reinforcement Learning from Human Feedback (RLHF) literature [36, 55, 51, 2, 3], we choose REINFORCE [54] ($\text{DDPO}_{\text{SF}}$) over PPO [45] ($\text{DDPO}_{\text{IS}}$) for its superior training stability. We provide two hypotheses behind our surprising finding.
@@
(2) The RL Finetuning (RLFT) sample efficiency is less important for a large model which requires less finetuning steps, as demonstrated in studies of LLM instruction finetuning [10]. Similarly, our RLFT on a 2.6B-parameter UNet from SDXL [39] only takes 55 epochs, as opposed to DDPO’s [5] RLFT on a 860M-parameter UNet from SD 1.4 [42] using 200 epochs. Therefore, the potential sample efficiency gain provided by the multi-step update of PPO [45] gets outweighted by the training stability provided by REINFORCE [54].
@@ C.1 Training Data
An advantage of Reinforcement Learning Finetuning (RLFT) over Supervised Finetuning (SFT) is that, we can manually create a high-quality text prompts training set, while creating a dataset of diverse ground truth multi-view images for these high-quality text prompts is prohibitively expensive for SFT. By relying on samples generated by the model itself to compute the reward and the loss, RLFT can optimize a model beyond the limitation of a dataset and preserves the diversity and the style of the base model. In Carve3D, our training prompts preparation process involves two strategies.
@@
Instead of randomly sampling prompts from a dataset, we employ a data curation strategy where prompts with lowest rewards are selected. Specifically, we run inference of the base model on a prompt dataset, generating four results per prompt, compute the MRC rewards for each result, and sort the prompts according to their average reward. This is derived from observation that, for certain prompts, the model generates nearly optimal outputs with rewards close to the rewards of ground truth views of a 3D asset [12] (Fig. 4). Thus, the curated lowest-reward prompts have substantial room for 3D consistency improvement and prevent learning stagnation. This approach not only brings more efficient training but also provides a more generalized improvement in 3D consistency to the testing set.
@@
The prompt dataset from DreamFusion [40], which contains 414 prompts and is commonly used as testing set. To employ the DreamFusion prompt set also as our testing set, we create a new prompt dataset with ChatGPT4 [35]. Following our training data curation strategy, we first sort the DreamFusion [40] prompts according to their rewards attained by the base Instant3D [24] model. We provide the sorted prompt set to ChatGPT4, and ask it to summarize the characteristics of the low-reward prompts by looking at the low-reward, median-reward, and high-reward prompts. ChatGPT4 summarizes low-reward prompts to possess properties of “complex and creative”. We then ask it to generate 100 low-reward prompts that are both complex and creative, and another 100 low-reward prompts that are “complex but not too creative". For each set, again, we sort the prompts according to their rewards, and select those with the lowest rewards to be our training prompt set. Our best results are obtained with the “complex but not too creative" set.
@@ C.2 Training Details
We use minibatches of size 8 during sampling and 4 during training due to the limited GPU memory. The total batch size of 768 is evenly distributed among each GPU, so that the per GPU batch size is 16. The model samples two minibatches of size 8 on all GPUs to reach the total batch size. Similarly, the model accumulates gradients over four minibatches of size 4 on all GPUs, before synchronizing the gradients and performing an optimizer step. We use a per prompt stat tracker with windows of size 76, so that it roughly tracks all the rewards per prompt ever 3 epochs. This is much larger than DDPO’s default tracking window of size 32 for better training stability. The coefficients for the advantage terms in Eq. 10 are $\alpha=1$ and $\beta=0.2$.
@@
C.3 Experiment Details
All quantitative and qualitative experiments and the user study uses the results from evaluating each model on the DreamFusion [40] testing prompt set, containing 415 prompts, for 4 times, which equals to 1660 results per model.
For fair qualitative comparisons, results from each model are generated from the the same initial noise. Since Instant3D-20K and -100K and Carve3D are all finetuned from Instant3D-10K, their output tend to represent the same object when given the same initial noise (e.g. Figs. 1, 6 and 11).
Since Zero123++ [47] and SyncDreamer [28] are image-to-multi-view diffusion models, we let them take one of Carve3D’s output image and their input image-conditioning. Therefore, their output has the same level of prompt alignment, texture details, and diversity as Carve3D.
@@ Appendix D Additional MRC Metric Experiments
Here, we show the full results for the metric experiments for the inpainting distortion (Fig. 14) discussed in Secs. 3.3, 3 and 4. We also conduct metric experiments with other distortions types: azimuth rotation (Fig. 15, and elevation rotation (Fig. 16). In azimuth and elevation rotation, for one out of the four views, we rotate the object with an azimuth or elevation rotation by $3.6$ or $4$ degrees, before rendering that view, and also use the original camera extrinsic matrix as the input to Sparse View LRM [24, 17]. The quantitative results matches our expectations, where MRC, i.e. with LPIPS, monotonically decreases as we intentionally add more distortion.
@@
Here, we compare substituting LPIPS [57] with L1, L2, PSNR, and SSIM in the Multi-view Reconstruction Consistency (MRC) metric experiments on all distortion types. In the inpainting distortion experiments (Fig. 14), which is the most representative of diffusion model’s inconsistencies, LPIPS is more linear than other pixel level image metrics. In azimuth and elevation distortion experiments (Figs. 15 and 16), all image metrics shows monotonically decreasing pattern, while pixel-level image metrics are more linear. This is expected as the distortion is pixel-aligned and more structured.
Figure 9: Alation study on the boundingbox normalization of LPIPS for MRC evaluation. Top: with boundingbox normalization, the size of the foreground object is similar that of the base model. Bottom: without boundingbox normalization, the size of the foreground object after RL finetuning is substantially smaller than that of the base model.
Figure 10: Ablation study on KL divergence regularization. Top: KL Divergence between the base model and the finetuned model on testing set. Bottom: mean MRC reward on testing set. Our KL divergence regularization does not sacrifice the model’s sample efficiency of the reward curve. The finetuned model has a lower KL divergence to the base model when using KL divergence regularization than without, which prevents degraded object identity and reduced texture details.